Express the following surfaces in spherical coordinates.

i.xz=3

ii.x²+y²-z²=1


1
Expert's answer
2021-02-04T07:54:29-0500

"\begin{cases}
 x=rsin\theta cos \psi ,\\
y=rsin\theta sin \psi ,\\
z=rcos\theta .
\end{cases}"

1.

"r^2sin\theta cos \theta cos \psi=3."

2.

"r^2cos2\theta=-1."